# A Smooth Approximation Framework for Weakly Convex Optimization

### v1

@@ 3.2 Partial smooth minimization of composite problems
where only $f$ is replaced by its smooth approximation $f_{\eta}$. We assume that the smooth approximation $f_{\eta}$ is computationally tractable, namely, its (approximate) first-order information is available via certain oracle calls. To analyze algorithms for solving (13), we consider two commonly used stationarity measures from the literature. The first is the norm of the generalized gradient, which is closely associated with the proximal gradient method [Saeed2016accelerated, nesterov2013gradient]. Given a stepsize $\gamma>0$, the generalized gradient is defined by:
@@
When $\mathcal{G}_{\gamma}(\mathbf{x})=\mathbf{0}$, $\mathbf{x}$ satisfies the first-order stationarity condition for (2). Alternatively, davis2019stochastic propose the gradient norm of the Moreau envelope: $\lVert\nabla\phi_{\eta}^{\beta}(\mathbf{x})\rVert$ as a stationarity measure. The quantitative relationship between these two measures has been explored by drusvyatskiy2019efficiency. The following theorem connects these criteria for the smoothed problem to the approximate stationarity of the original nonsmooth problem.
@@
Let $\mathbf{x}\in\operatornamewithlimits{dom}\phi$ and $f_{\eta}$ be a $(\bar{\rho},\eta)$-smooth approximation of a $\rho$-weakly convex function $f$.
@@
    If $\lVert\nabla\phi_{\eta}^{\beta}(\mathbf{x})\rVert\leq\varepsilon$, then $\mathbf{x}$ is a $\Big(\sqrt{\frac{2\mathcal{R}_{\eta}(\mathbf{y}_{\mathbf{x}})}{\beta-\rho}}+\beta^{-1}\varepsilon,\beta\sqrt{\frac{2\mathcal{R}_{\eta}(\mathbf{y}_{\mathbf{x}})}{\beta-\rho}}+\varepsilon\Big)$-stationary point of problem (1), where $\mathbf{y}_{\mathbf{x}}=\mathrm{prox}_{\phi_{\eta}/\beta}(\mathbf{x})$.
@@ 4.1 Generalized Nesterov’s smooth approximation
Building on the seminal work of nesterov2005smooth, we study nonsmooth functions that can be expressed as the composition of a convex function with a nonlinear map:
@@
Let $h^{*}(\mathbf{y})=\sup_{\mathbf{z}\in\mathbb{V}}\{\langle\mathbf{z},\mathbf{y}\rangle-h(\mathbf{z})\}$ be the convex conjugate of $h$. Due to bi-conjugacy (e.g. [borwein2006convex, Theorem 4.2.1]), we can express $f$ as
@@
Part 3). Lipschitz smoothness of $h_{\eta}$ follows from the Baillon-Haddad theorem [Beck2012smoothing, Lemma 4.1]. ∎
@@
$\quad+\nabla h_{\eta}(A(\mathbf{x}_{2}))^{\top}\big[A(\mathbf{x}_{1})-A(\mathbf{x}_{2})-\nabla A(\mathbf{x}_{2})(\mathbf{x}_{1}-\mathbf{x}_{2})\big]$ (21)
@@
$\nabla h_{\eta}(A(\mathbf{x}_{2}))^{\top}\big[A(\mathbf{x}_{1})-A(\mathbf{x}_{2})-\nabla A(\mathbf{x}_{2})(\mathbf{x}_{1}-\mathbf{x}_{2})\big]$
@@
$\geq-\frac{BL_{A}}{2}\lVert\mathbf{x}_{1}-\mathbf{x}_{2}\rVert^{2}.$ (23)
Combining (21) and (23) gives the desired weak convexity.
@@
    Bounded Domain: If the domain $\mathcal{X}$ is compact, then the continuous map $\nabla A$ is bounded, i.e., $M\coloneqq\sup_{\mathbf{x}\in\mathcal{X}}\|\nabla A(\mathbf{x})\|_{\mathrm{op}}<\infty$. Consequently, $\mathcal{L}_{\eta}$ can be uniformly bounded as: $\mathcal{L}_{\eta}(\mathbf{x}_{1},\mathbf{x}_{2})\leq\frac{M^{2}}{\sigma\eta}$.
@@
Consider the following max-type function
@@
where each $f_{j}(\mathbf{x})$ is smooth. This is a special case of (18) where $h(\mathbf{z})=\max_{1\leq j\leq m}y_{i}=\max_{\mathbf{y}\in\mathcal{S}^{m}}\langle\mathbf{y},\mathbf{z}\rangle$ and $\mathcal{S}^{m}=\{\mathbf{y}\in\mathbb{R}^{m}:\langle\mathbf{1}_{m},\mathbf{y}\rangle=1,\mathbf{y}\geq\mathbf{0}\}$ is the probability simplex. Define the prox-function $\omega(\mathbf{y})=\sum_{i=1}^{m}y_{i}\log y_{i}$ when $\mathbf{y}\in\mathcal{S}^{m}$ and $\omega(\mathbf{y})=+\infty$ otherwise. Then the smoothed approximation is given by
$f_{\eta}(\mathbf{x})=\max_{\mathbf{y}\in\mathcal{S}^{m}}\{\langle\mathbf{y},F(\mathbf{x})\rangle-\eta\omega(\mathbf{y})\}=\eta\log\Big(\sum_{i=1}^{m}\exp\Big(\frac{f_{j}(\mathbf{x})}{\eta}\Big)\Big),$ (25)
@@ 4.2 Moreau envelope smoothing
We present several properties regarding the weak convexity and Lipschitz smoothness of the Moreau envelope. While some of these results appear in prior works [davis2019stochastic, hoheisel2010proximal], we include them here for completeness.
@@
    For any $\mathbf{x}\in\operatornamewithlimits{dom}f$, we have
    $f^{\beta}(\mathbf{x})+\frac{(1-\rho/\beta)}{2\beta}\lVert\nabla f^{\beta}(\mathbf{x})\rVert^{2}\leq f(\mathbf{x})\leq f^{\beta}(\mathbf{x})+\frac{\lVert\partial f(\mathbf{x})\rVert^{2}}{\beta-\rho}.$
@@
$\mathcal{R}_{\eta}(\mathbf{x})=\frac{\eta}{\max\{1,\rho\eta\}}\lVert\partial f(\mathbf{x})\rVert^{2},\quad\mathcal{L}_{\eta}(\mathbf{x},\mathbf{y})=2\rho+\eta^{-1}.$ (27)
@@
In view of Part 2) of Proposition 4.2, we have
$f^{\beta}(\mathbf{x})-f^{\beta}(\mathbf{y})+\langle\nabla f^{\beta}(\mathbf{y}),\mathbf{x}-\mathbf{y}\rangle$ $\geq-\frac{\rho}{2(1-\rho/\beta)}\lVert\mathbf{y}-\mathbf{x}\rVert^{2}$
@@
Hence $f_{\eta}$ is $2\rho$-weakly convex. Furthermore, Part 2) of Proposition 4.2 implies
@@
Part 3) of Proposition 4.2 implies $f_{\eta}(\mathbf{x})\leq f(\mathbf{x})\leq{\lVert\partial f(\mathbf{x})\rVert^{2}}/{\max\{\eta^{-1},\rho\}}$. ∎
@@
Theorem 4.2 implies that for a fixed $\mathbf{x}$, the error term $\mathcal{R}_{\eta}(\mathbf{x})$ is $\mathcal{O}(\min\{\eta,\rho^{-1}\})$, while the smoothness constant of the approximation is $\mathcal{O}(\eta^{-1})$. The error term differs slightly from that of convex smoothing, where $\mathcal{R}_{\eta}(\mathbf{x})=R\eta$ for some $R>0$ [beck2017first]. The difference arises as applying the Moreau envelope to a nonconvex function can increase its negative curvature (i.e., yield a larger weak convexity constant), as shown in (26) of Proposition 4.2.
@@
for a given reference point $\hat{\mathbf{x}}\in\mathbb{R}^{d}$. This problem can sometimes be solved in closed form [davis2019stochastic, Section 5], or more generally via the prox-linear algorithm
@@
which admits linear convergence guarantees [deng2021minibatch, Section D.3]. When $m$ is moderate and $h$ has a simple structure (e.g., $h(\cdot)=|\cdot|$), it is advantageous to reformulate the problem as a min-max saddle-point problem:
@@
We now compare the generalized Nesterov smoothing with the classical Moreau envelope smoothing for the composite function $h(A(\mathbf{x}))$. The two techniques are closely related and are, in certain cases, identical. Specifically, suppose $A$ is the identity map $A(\mathbf{x})=\mathbf{x}$, then we have $h^{*}(\mathbf{y})=f^{*}(\mathbf{y})$. In this case, Nesterov’s smoothing reduces to the infimal convolution smooth approximation. If we choose the prox-function to be the squared norm, $\omega(\mathbf{y})=\frac{1}{2}\|\mathbf{y}\|^{2}$, then it can be further interpreted as the dual formulation of Moreau envelope smoothing. See [beck2017first, Section 4.4]. For a general nonlinear map, the two smoothing approaches are different, revealing a fundamental trade-off:
@@
    Nesterov’s smoothing is often more computationally tractable since the subproblem in (20) is solved over the dual space $\operatornamewithlimits{dom}h^{*}$. It can be more efficient if the structure of $h^{*}$ is simple. In contrast, computing the Moreau envelope of $f$ requires evaluating a nontrivial proximal operator, which often requires subroutines. An exception is in stochastic optimization, where the proximal subproblem involving only a single sample sometimes admits a closed-form solution [davis2019stochastic, Section 5].
@@ 4.3.2 Expectation function
Applying [ruszczynski2003optimality, Proposition 2], we can show $f_{\eta}(\cdot,\xi)$ is locally Lipschitz and differentiable:
@@ 4.4 More composite functions
We now examine objective functions of the form $f(\mathbf{x})=g(A\mathbf{x})$, where $g$ is a $\rho$-weakly convex function and $A$ is a linear operator. This class of problems is studied in bohm2021variable.
@@
$\tilde{\rho}=\bar{\rho}\lVert A\rVert_{\mathrm{op}}^{2},\quad\tilde{R}_{\eta}(\mathbf{x})=\mathcal{R}_{\eta}(A\mathbf{x}),\quad\tilde{\mathcal{L}}_{\eta}(\mathbf{x},\mathbf{y})=\lVert A\rVert_{\mathrm{op}}L_{\eta}(A\mathbf{x},A\mathbf{y}).$
@@ 5 Smoothing algorithms
Condition (34) is satisfied, for example, when $\operatornamewithlimits{dom}r$ is a compact subset in the interior of $\mathcal{X}$; see Remark 4.3 and Remark 4.1 for further discussions. We restrict $\eta$ to a bounded interval since, in general, $\mathcal{R}_{\eta}$ may not be linear with respect to $\eta$ over $\mathbb{R}$ (as occurs, for instance, with Moreau envelope smoothing). This restriction to $\eta\in(0,\theta]$ is mild, since our analysis ultimately adopts a choice of $\eta=\mathcal{O}(\varepsilon^{2})$. Note that while we assume $\mathcal{R}_{\eta}$ and $\mathcal{L}_{\eta}$ are independent of the variables $\mathbf{x}$ and $\mathbf{y}$, this assumption will be relaxed to more general settings in which these quantities depend on the variables, as discussed in Section 6. The following assumption regarding the bounded variance of the gradient is standard in the literature [lan2020first].
@@ 3 $\mathbf{x}^{k}=\mathrm{prox}_{\gamma r}(\mathbf{x}^{k-1}-\gamma\mathbf{g}^{k-1})$;
There are two ways to analyze Algorithm 1. One follows the analysis in ghadimi2016mini, which can be interpreted as a perturbed proximal gradient method. The batch size $m$ needs to be chosen sufficiently large to control the stochastic approximation error.
@@
When $\sigma=0$, Algorithm 1 reduces to the proximal gradient method. In order to obtain an $(\varepsilon,\varepsilon)$-stationary point of (1), according to Theorem 3.2, we would need $\eta=\mathcal{O}(\varepsilon^{2})$. Under this setting, Theorem 5.1 establishes an $\mathcal{O}(1/\varepsilon^{4})$ complexity bound. This complexity matches, but does not improve upon, the worst-case complexity of the proximal subgradient method [davis2019stochastic]. The lack of improvement is attributed to the poor conditioning (large $L_{\eta}$) introduced by smoothing.
@@
This complexity is even inferior to the $\mathcal{O}(1/\varepsilon^{4})$ result of the standard stochastic subgradient method. Fortunately, this suboptimal bound is an artifact of the analysis. A sharper complexity can be derived by analyzing the algorithm through the lens of the Moreau envelope [davis2019stochastic], which allows an arbitrary batch size. Following the notation from the previous sections, we denote the Moreau envelope of $\phi_{\eta}$ by
@@
By leveraging the Moreau envelope as a potential function and following the analysis of davis2019stochastic, we establish the following convergence guarantee.
@@
Suppose $\Delta\geq\phi(\mathbf{x}^{0})-\min_{\mathbf{x}}\phi(\mathbf{x})$. If we take $m=1$, $\gamma=(c\sqrt{K}+\hat{\rho}+L_{\eta})^{-1}$ where $c=\sqrt{\frac{\hat{\rho}}{(\Delta+R\eta)}}\sigma$, then
$\min_{0\leq k\leq K-1}\mathbb{E}[\lVert\nabla\phi_{\eta}^{\hat{\rho}}(\mathbf{x}^{k})\rVert^{2}]\leq\frac{\hat{\rho}}{\hat{\rho}-\bar{\rho}}\bigg\{\frac{(2\hat{\rho}-\bar{\rho}+L_{\eta})(\Delta+R\eta)}{K}+\sqrt{\frac{\hat{\rho}(\Delta+R\eta)}{K}}\sigma\bigg\}.$ (35)
@@
which matches the sample complexity bound achieved by the stochastic subgradient method [davis2019stochastic].

### v3

@@ 3.2 Partial smooth minimization of composite problems
where only $f$ is replaced by its smooth approximation $f_{\eta}$. We assume that the smooth approximation $f_{\eta}$ is computationally tractable, namely, its (approximate) first-order information is available via certain oracle calls. To analyze algorithms for solving (12), we consider two commonly used stationarity measures from the literature. The first is the norm of the generalized gradient, which is closely associated with the proximal gradient method [22, 35]. Given a stepsize $\gamma>0$, the generalized gradient is defined by:
@@
When $\mathcal{G}_{\gamma}(\mathbf{x})=\mathbf{0}$, $\mathbf{x}$ satisfies the first-order stationarity condition for (2). Alternatively, Davis and Drusvyatskiy [13] propose the gradient norm of the Moreau envelope: $\lVert\nabla\phi_{\eta}^{\beta}(\mathbf{x})\rVert$ as a stationarity measure. The quantitative relationship between these two measures has been explored by Drusvyatskiy and Paquette [16]. The following theorem connects these criteria for the smoothed problem to the approximate stationarity of the original nonsmooth problem.
@@
Suppose that $f$ is $\rho$-weakly convex and $f_{\eta}$ is a $(\bar{\rho},\eta)$-smooth approximation of $f$. Let $\mathbf{x}\in\operatornamewithlimits{dom}\phi$.
@@
    Assume $\beta>\max\{\rho,\bar{\rho}\}$. If $\lVert\nabla\phi_{\eta}^{\beta}(\mathbf{x})\rVert\leq\varepsilon$, then $\mathbf{x}$ is a $\Big(\sqrt{\frac{2\mathcal{R}_{\eta}(\mathbf{y}_{\mathbf{x}})}{\beta-\rho}}+\beta^{-1}\varepsilon,\beta\sqrt{\frac{2\mathcal{R}_{\eta}(\mathbf{y}_{\mathbf{x}})}{\beta-\rho}}+\varepsilon\Big)$-stationary point of problem (1), where $\mathbf{y}_{\mathbf{x}}=\mathrm{prox}_{\phi_{\eta}/\beta}(\mathbf{x})$.
@@ 4.1 Generalized Nesterov’s smooth approximation
Building on the seminal work of Nesterov [34], we study nonsmooth functions that can be expressed as the composition of a convex function with a nonlinear map:
@@
Let $h^{*}(\mathbf{y})=\sup_{\mathbf{z}\in\mathbb{V}}\{\langle\mathbf{z},\mathbf{y}\rangle-h(\mathbf{z})\}$ be the convex conjugate of $h$. Due to bi-conjugacy (e.g. [6, Theorem 4.2.1]), we can express $f$ as
@@
Part 3). Lipschitz smoothness of $h_{\eta}$ follows from the Baillon-Haddad theorem [2, Lemma 4.1]. ∎
@@
$\quad+\nabla h_{\eta}(A(\mathbf{x}_{2}))^{\top}\big[A(\mathbf{x}_{1})-A(\mathbf{x}_{2})-\nabla A(\mathbf{x}_{2})(\mathbf{x}_{1}-\mathbf{x}_{2})\big]$
@@
$\nabla h_{\eta}(A(\mathbf{x}_{2}))^{\top}\big[A(\mathbf{x}_{1})-A(\mathbf{x}_{2})-\nabla A(\mathbf{x}_{2})(\mathbf{x}_{1}-\mathbf{x}_{2})\big]$ (22)
@@
$\geq-\frac{BL_{A}}{2}\lVert\mathbf{x}_{1}-\mathbf{x}_{2}\rVert^{2}.$
Combining (20) and ($2$) gives the desired weak convexity.
@@
    Bounded Domain: If the domain $\mathcal{X}$ is compact, then the continuous map $\nabla A$ is bounded, i.e., $M\coloneqq\sup_{\mathbf{x}\in\mathcal{X}}\|\nabla A(\mathbf{x})\|_{\mathrm{op}}<\infty$. Hence, $\mathcal{L}_{\eta}$ can be uniformly bounded as: $\mathcal{L}_{\eta}(\mathbf{x}_{1},\mathbf{x}_{2})\leq\frac{M^{2}}{\sigma\eta}+BL_{A}$.
@@
Consider the max-type function
@@
where each $f_{j}(\mathbf{x})$ is smooth. This is a special case of (17) where $h(\mathbf{z})=\max_{1\leq j\leq m}z_{j}=\max_{\mathbf{y}\in\mathcal{S}^{m}}\langle\mathbf{y},\mathbf{z}\rangle$ and $\mathcal{S}^{m}=\{\mathbf{y}\in\mathbb{R}^{m}:\langle\mathbf{1}_{m},\mathbf{y}\rangle=1,\mathbf{y}\geq\mathbf{0}\}$ is the probability simplex. Define the prox-function $\omega(\mathbf{y})=\sum_{i=1}^{m}y_{i}\log y_{i}+\log m$ when $\mathbf{y}\in\mathcal{S}^{m}$ and $\omega(\mathbf{y})=+\infty$ otherwise. Then the smoothed approximation is given by
$f_{\eta}(\mathbf{x})=\max_{\mathbf{y}\in\mathcal{S}^{m}}\{\langle\mathbf{y},F(\mathbf{x})\rangle-\eta\omega(\mathbf{y})\}=\eta\log\Big(\sum_{i=1}^{m}\exp\Big(\frac{f_{i}(\mathbf{x})}{\eta}\Big)\Big)-\eta\log m,$ (24)
@@ 4.2 Moreau envelope smoothing
We present several properties regarding the weak convexity and Lipschitz smoothness of the Moreau envelope. While some of these results appear in prior works [13, 24], we include them here for completeness.
@@
    Let $\mathbf{x}\in\operatornamewithlimits{dom}f$ be subdifferentiable, we have
    $f^{\beta}(\mathbf{x})+\frac{(1-\rho/\beta)}{2\beta}\lVert\nabla f^{\beta}(\mathbf{x})\rVert^{2}\leq f(\mathbf{x})\leq f^{\beta}(\mathbf{x})+\frac{\lVert\partial f(\mathbf{x})\rVert^{2}}{2(\beta-\rho)}.$
@@
$\mathcal{R}_{\eta}(\mathbf{x})=\frac{\eta}{2\max\{1,\rho\eta\}}\lVert\partial f(\mathbf{x})\rVert^{2},\quad\mathcal{L}_{\eta}(\mathbf{x},\mathbf{y})=2\rho+\eta^{-1}.$ (26)
@@
In view of (25), we have
$f_{\eta}(\mathbf{x})-f_{\eta}(\mathbf{y})+\langle\nabla f_{\eta}(\mathbf{y}),\mathbf{x}-\mathbf{y}\rangle$ $\geq-\frac{\rho}{2(1-\rho/\beta)}\lVert\mathbf{y}-\mathbf{x}\rVert^{2}$
@@
Hence $f_{\eta}$ is $2\rho$-weakly convex. The choice of $\beta$ ensures $\beta\geq\rho/(1-\rho/\beta)$. Together with the Lipschitz continuity of $\nabla f^{\beta}$, we have
@@
Part 2) of Proposition 4.2 implies $f_{\eta}(\mathbf{x})\leq f(\mathbf{x})\leq{f_{\eta}(\mathbf{x})+\lVert\partial f(\mathbf{x})\rVert^{2}}/{2(\max\{\eta^{-1},\rho\})}$. ∎
@@
Theorem 4.2 implies that for a fixed $\mathbf{x}$, the error term $\mathcal{R}_{\eta}(\mathbf{x})$ is $\mathcal{O}(\min\{\eta,\rho^{-1}\})$, while the smoothness constant of the approximation is $\mathcal{O}(\rho+\eta^{-1})$. The error term differs slightly from that of convex smoothing, where $\mathcal{R}_{\eta}(\mathbf{x})=R\eta$ for some $R>0$ [3]. The difference arises as applying the Moreau envelope to a nonconvex function can increase its negative curvature (i.e., yield a larger weak convexity constant), as shown in (25) of Proposition 4.2.
@@
for a given reference point $\hat{\mathbf{x}}\in\mathbb{R}^{d}$. This problem can sometimes be solved in closed form [13, Section 5], or more generally via the prox-linear algorithm
@@
which admits linear convergence guarantees [15, Section D.3]. When $m$ is moderate and $h$ has a simple structure (e.g., $h(\cdot)=|\cdot|$), it is advantageous to reformulate the problem as a min-max saddle-point problem:
@@
We now compare the generalized Nesterov smoothing with the classical Moreau envelope smoothing for the composite function $h(A(\mathbf{x}))$. The two techniques are closely related and are, in certain cases, identical. Specifically, suppose $A$ is the identity map $A(\mathbf{x})=\mathbf{x}$, then we have $h^{*}(\mathbf{y})=f^{*}(\mathbf{y})$. In this case, Nesterov’s smoothing reduces to the infimal convolution smooth approximation. If we choose the prox-function to be the squared norm, $\omega(\mathbf{y})=\frac{1}{2}\|\mathbf{y}\|^{2}$, then it can be further interpreted as the dual formulation of Moreau envelope smoothing. See [3, Sec. 4.4]. For a general nonlinear map, the two smoothing approaches are different, revealing a fundamental trade-off:
@@
    Nesterov’s smoothing is often more computationally tractable since the subproblem in (19) is solved over the dual space $\operatornamewithlimits{dom}h^{*}$. It can be more efficient if the structure of $h^{*}$ is simple. In contrast, computing the Moreau envelope of $f$ requires evaluating a nontrivial proximal operator, which often requires subroutines. An exception is in stochastic optimization, where the proximal subproblem involving only a single sample sometimes admits a closed-form solution [13, Sec. 5].
@@ 4.3.2 Expectation function
Applying [37, Proposition 2], we can show $f_{\eta}(\cdot,\xi)$ is locally Lipschitz and differentiable:
@@ 4.4 More composite functions
We now examine objective functions of the form $f(\mathbf{x})=g(A\mathbf{x})$, where $g$ is a $\rho$-weakly convex function and $A$ is a linear operator. This class of problems is studied in Böhm and Wright [4].
@@
$\tilde{\rho}=\bar{\rho}\lVert A\rVert_{\mathrm{op}}^{2},\quad\tilde{R}_{\eta}(\mathbf{x})=\mathcal{R}_{\eta}(A\mathbf{x}),\quad\tilde{\mathcal{L}}_{\eta}(\mathbf{x},\mathbf{y})=\lVert A\rVert_{\mathrm{op}}\mathcal{L}_{\eta}(A\mathbf{x},A\mathbf{y}).$
@@ 5 Algorithms for smooth approximation
Condition (33) is satisfied, for example, when $\operatornamewithlimits{dom}r$ is a compact subset in the interior of $\mathcal{X}$; see Remark 4.3 and Remark 4.1 for further discussions. We restrict $\eta$ to a bounded interval since, in general, $\mathcal{R}_{\eta}$ may not be linear with respect to $\eta$ over $\mathbb{R}$ (as occurs, for instance, with Moreau envelope smoothing). This restriction to $\eta\in(0,\theta]$ is mild, since our analysis ultimately adopts a choice of $\eta=\mathcal{O}(\varepsilon^{2})$. Note that while we assume $\mathcal{R}_{\eta}$ and $\mathcal{L}_{\eta}$ are independent of the variables $\mathbf{x}$ and $\mathbf{y}$, this assumption will be relaxed to more general settings in which these quantities depend on the variables, as discussed in Section 6. The following assumption regarding the bounded variance of the gradient is standard in the literature [27].
@@ 3 $\mathbf{x}^{k}=\mathrm{prox}_{\gamma r}(\mathbf{x}^{k-1}-\gamma\mathbf{g}^{k-1})$;
There are two ways to analyze Algorithm 1. One follows the analysis in Ghadimi et al. [20], which can be interpreted as a perturbed proximal gradient method. The batch size $m$ needs to be chosen sufficiently large to control the stochastic approximation error.
@@
When $\sigma=0$, Algorithm 1 reduces to the proximal gradient method. In order to obtain an $(\varepsilon,\varepsilon)$-stationary point of (1), according to Theorem 3.2, we would need $\eta=\mathcal{O}(\varepsilon^{2})$. Under this setting, Theorem 5.1 establishes an $\mathcal{O}(1/\varepsilon^{4})$ complexity bound. This complexity matches, but does not improve upon, the worst-case complexity of the proximal subgradient method [13]. The lack of improvement is attributed to the poor conditioning (large $L_{\eta}$) introduced by smoothing.
@@
This complexity is even inferior to the $\mathcal{O}(1/\varepsilon^{4})$ result of the standard stochastic subgradient method. Fortunately, this suboptimal bound is an artifact of the analysis. A sharper complexity can be derived by analyzing the algorithm through the lens of the Moreau envelope [13], which allows an arbitrary batch size. Following the notation from the previous sections, we denote the Moreau envelope of $\phi_{\eta}$ by
@@
By leveraging the Moreau envelope as a potential function and following the analysis of Davis and Drusvyatskiy [13], we establish the following convergence guarantee.
@@
Suppose $\Delta\geq\phi(\mathbf{x}^{0})-\min_{\mathbf{x}}\phi(\mathbf{x})$. Let $m=1$, $\gamma=(c\sqrt{K}+\hat{\rho}+L_{\eta})^{-1}$ where $c=\sqrt{\frac{\hat{\rho}}{2(\Delta+R\eta)}}\sigma$, then
$\min_{0\leq k\leq K-1}\mathbb{E}[\lVert\nabla\phi_{\eta}^{\hat{\rho}}(\mathbf{x}^{k})\rVert^{2}]\leq\frac{\hat{\rho}}{\hat{\rho}-\bar{\rho}}\Big\{\frac{(2\hat{\rho}-\bar{\rho}+L_{\eta})(\Delta+R\eta)}{K}+\sqrt{\frac{2\hat{\rho}(\Delta+R\eta)}{K}}\sigma\Big\}.$ (34)
@@
which matches the sample complexity bound achieved by the stochastic subgradient method [13].